Andy and Chris work together at a bakery making lemonade and iced tea. Andy can make 11 cups of lemonade per hour or 20 cups of iced tea per hour or some combination of both. Chris can make 14 cups of lemonade per hour or 30 cups of iced tea per hour or some combination of both. Their productivity can be summarized in the following table:

Lemonade Iced Tea
Andy 11 20
Chris 14 30

Fill in the blanks for the statements below. If your answer is a number, put in the form of a decimal. For example if your answer is 4/3 you should put 1.33. If your answer is a name make sure to spell it exactly as it's spelled above.

Andy's opportunity cost of 1 cup of lemonade is  cups of iced tea. Andy's opportunity cost of 1 cup of iced tea is  cups of lemonade.

Chris's opportunity cost of 1 cup of lemonade is  cups of iced tea. Chris's opportunity cost of 1 cup of iced tea is  cups of lemonade .

has the absolute advantage in making lemonade .  has the absolute advantage in making iced tea .

has the comparative advantage in making lemonade .  has the comparative advantage in making iced tea .

Suppose that Andy and Chris are not specialized. Andy is making 3.85 cups of lemonade AND 13 cups of iced tea per hour. Chris is making 6.53 cups of lemonade AND 16 cups of iced tea per hour.

If Andy and Chris fully specialize they will all together make  more cups of lemonade per hour and  more cups of iced tea per hour.
Answer #1

Andy:

Andy can make a maximum of 11 cups of lemonade per hour or 20 cups of iced tea per hour

Andy's opportunity cost of 1 cup of lemonade = 20/11 = 1.82 cups of iced tea (1)

Andy's opportunity cost of 1 cup of iced tea = 11/20 = 0.55 cups of lemonade (2)

Chris:

Chris can make a maximum of 14 cups of lemonade per hour or 30 cups of iced tea per hour

Chris's opportunity cost of 1 cup of lemonade = 30/14 = 2.14 cups of iced tea (3)

Chris's opportunity cost of 1 cup of iced tea = 14/30 = 0.47 cups of lemonade (4)

Absolute Advantage:

It can be observed that Chris can produce a higher quantity of both lemonade & iced tea in an hour. Therefore, Chris(5) has the absolute advantage in making lemonade. Chris(6) has the absolute advantage in making iced tea

Comparative Advantage:

The opportunity cost of making lemonade is lower for Andy. Therefore, Andy(7) has the comparative advantage in making lemonade.

The opportunity cost of making iced tea is lower for Chris. Therefore, Chris(8) has the comparative advantage in making iced tea.

Before and after specialization:

Before -

Andy is making 3.85 cups of lemonade AND 13 cups of iced tea per hour. Chris is making 6.53 cups of lemonade AND 16 cups of iced tea per hour.

Total production per hour:

Lemonade: 3.85 + 6.53 = 10.38 cups

Iced tea: 13 + 16 = 29 cups

After -

Andy specializes in making lemonade and Chris specializes in making iced tea.

Andy makes 11 cups of lemonade and Chris makes 30 cups of iced tea

Total production:

Lemonade: 11 cups

Iced tea: 30 cups

Gains from specialization:

Lemonade: 11 - 10.38 = 0.62 cups (9)

Iced tea: 30 - 29 = 1 cup (10)